how many ounces of water must be added to 20 ounces of 8% salt solution to make 2% salt soltion
Answer by gonzo(654)

20 ounces of 8% salt solution yields 1.6 ounces of salt and 18.4 ounces of water.
if you want 1.6 ounces of salt to be 2% of a solution, then you divide 1.6 by .02 to get 80 ounces of solution. 1.6 / 80 equates to .02 so this new solution is 1.6 ounces of salt and 78.4 ounces of water.
since you already have 18.4 ounces of water in the original solution, you need to add (78.4-18.4) = 60 ounces of water.
you take your original solution of 20 ounces and add 60 ounces of water.
the original solution had 1.6 ounces of salt and 18.4 ounces of water to make a total of 20 ounces of solution.
the new solution has 1.6 ounces of salt and 78.4 ounces of water to make a total of 80 ounces of solution.
1.6 / 20 = .08
1.6 / 80 = .02